Assume the terminal value of an investment is $27,500. This purchase has a cost of $50,000 with a marginal tax rate of 13%. This investment has an annual depreciation of $4,000. It is required at least a 10% pretax rate of return on capital. What is the present value of after-tax terminal value after 4 years?
Answer #1

value after 4 years = 50,000 - 4,000*4 = 34,000

loss = 34,000 - 27,500 = 6,500

tax advantage = 013 * 6500 = 845

after-tax terminal value = 27,500 + 845 = 28,345

PV of after tax terminal value = 28,345/1.10^4 = 19,360.02